0 0
Add a comment Improve this question Transcribed image text
Answer #1

a) it is a simple lo pass rc filter

b) fc = 2/2pirc

fc = 1000/2*3.14*1*0.1

fc = 1592.4 Hz

c) fc = 1/2pisqrt(r1c1r2c2)